Hi. I have been having difficulty with word problems. I am in algebra 1. The problem on my paper says, "Nate's father is 25 years older than Nate. Nate's father is 7 years older than twice Nate's age. How old is Nate's father?" If you could talk me through it, that would be great. Thank you.
Answer by stanbon(75887)

"Nate's father is 25 years older than Nate. Nate's father is 7 years older than twice Nate's age. How old is Nate's father?"
-------------------------
Equations:
f = n + 25
f = 2n+7
-----
Solve by substitution::
n+25 = 2n +7
n = 15 (Nate's age now)
----
Solve for "f":
f = n+25 = 15+25 = 40 (Father's age now)
-----------------
Cheers,
Stan H.
=================